### v1

# TAT-LLM: A Specialized Language Model for Discrete Reasoning over Tabular and Textual Data

Abstract.
In this work, we address question answering (QA) over a hybrid of tabular and textual data that are very common content on the Web (e.g. SEC filings), where discrete reasoning capabilities are often required. Recently, large language models (LLMs) like GPT-4 have demonstrated strong multi-step reasoning capabilities. We then consider harnessing the amazing power of LLMs to solve our task. We abstract a Step-wise Pipeline for tabular and textual QA, which consists of three key steps, including Extractor, Reasoner and Executor, and initially design an instruction to instantiate the pipeline and validate that GPT-4 outperforms all existing methods. However, utilizing an online LLM like GPT-4 holds various challenges in terms of cost, latency, and data security risk, which motivates us to specialize smaller LLMs in this task. We develop a TAT-LLM language model by fine-tuning LLaMA 2 with the training data generated automatically from existing expert-annotated datasets following the Step-wise Pipeline. The experimental results have verified that our TAT-LLM model can outperform all baseline models, including the previous best fine-tuned models and very large-scale LLMs like GPT-4 on FinQA, TAT-QA and TAT-DQA benchmarks. We hope our work can serve as a pioneering example of specializing smaller language models for specific tasks.
Question Answering, Large Language Model, Tabular and Textual Data, Discrete Reasoning, Natural Language Processing
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Question answeringccs: Computing methodologies Natural language processing
1. Introduction
Figure 1. Examples of QA with discrete reasoning over a hybrid of tabular and textual data.
Nowadays the Web inundates its users with vast and intricate content, necessitating content analysis techniques. In this work, we focus on analyzing a prevalent category of Web content: documents containing tables and text, including but not limited to SEC filings, financial statements, academic papers, and medical reports. These documents often feature extensive numerical data in both tabular and textual formats, demanding discrete reasoning for intelligent comprehension. Recent research (Zhu et al., 2021; Chen et al., 2021; Zhu et al., 2022) investigates the intelligent comprehension of such documents through question answering (QA) tasks, as exemplified in Figure 1. The model, provided with a table and relevant text as the context, needs to perform various types of discrete reasoning, such as arithmetic calculations, making comparisons, and counting, to answer the question.
To perform QA over hybrid tabular and textual data, a straightforward approach (Ran et al., 2019; Lei et al., 2022) involves taking the table, text, and question as input and generating the answer directly. However, this approach can be ineffective due to the complex reasoning process involved (Wei et al., 2022c). To address this, some works (Chen et al., 2021; Zhu et al., 2023a; Zhou et al., 2022; Li et al., 2023) decompose the task into multiple steps, producing intermediate results that serve as references for the final answer. For example, SoarGraph (Zhu et al., 2023a) first extracts the relevant information by selecting nodes in a constructed table-text graph structure and then employs a tree-based method to explicitly generate the mathematical expression of the answer. These multi-step approaches typically require distinct modules for each step and often optimize them concurrently through multi-task learning. However, the problem of how to decompose and model the answer process remains an open challenge.
We identify three key steps in the process of tabular and textual QA, and abstract a Step-wise Pipeline, as illustrated in Figure 2 b). In particular, 1) Extractor identifies the relevant information or evidence to the question from the given context; 2) Reasoner generates a mathematical equation or logic rule with the obtained information; and 3) Executor derives the final answer by executing the mathematical equation or logic rule with the associated information. These three steps emphasize different abilities for tabular and textual QA — understanding of the question and context, knowledge (e.g., logic) of answering the question, and precision in calculating the answer. As these steps produce a sequence of intermediate results, we can specifically model them and enhance their discrete reasoning ability given a practical scenario of tabular and textual QA problems.
A large language model (LLM) like GPT-3 (Brown et al., 2020) or FLAN (Wei et al., 2022a) has exhibited strong abilities required in the steps of the Step-wise Pipeline and an emergent multi-step reasoning ability (Wei et al., 2022b) with proper instructions such as chain-of-thought (CoT) (Wei et al., 2022c) and least-to-most (Zhou et al., 2023). We initially design a task instruction to instantiate the pipeline and validate that GPT-4 (OpenAI, 2023) outperforms existing methods for such QA tasks on multiple benchmarks (see Tables 24). However, utilizing an online LLM presents challenges in terms of cost, latency, and data security risks. On the other hand, fine-tuning a smaller language model specifically for math word problems has shown promising results (Cobbe et al., 2021a; Fu et al., 2023). This inspires us to explore the specialization of smaller language models for discrete reasoning over tabular and textual data within the Step-wise Pipeline.
To this end, we develop a TAT-LLM model by fine-tuning LLaMA 2 (Touvron et al., 2023b) with the training data generated automatically from existing expert-annotated datasets, as shown in Figure 2 c). In particular, we train the selected LLM to take the table, text, and question as input, and complete the task in three steps following the Step-wise Pipeline. We construct each of the training instances with five parts: Instruction, Table, Text, Question, and Response, from available tabular and textual QA datasets. It is worth mentioning that the model output (or response) is formatted to a structured table, with each row corresponding to one step of the pipeline, as shown in the right part of Figure 2 c). This design is friendly to further refining the final answer and automatic evaluation. Moreover, to enhance the model performance, we equip the TAT-LLM model with an External Executor, which strengthens the execution of logical rules and mathematical calculations to better infer the final answer. We test our TAT-LLM on three popular tabular and textual QA benchmarks: FinQA (Chen et al., 2021), TAT-QA (Zhu et al., 2021) and TAT-DQA (Zhu et al., 2022). The experimental results show that our smallest model TAT-LLM (7B) can outperform all baseline models and even beat GPT-4 on all three datasets.
We make the following main contributions in this work:
  • We abstract a Step-wise Pipeline including Extractor, Reasoner and Executor for better discrete reasoning capabilities and ground the pipeline on LLMs, showcasing their potential on the tabular and textual QA task.
  • We develop a TAT-GPT model by fine-tuning LLaMA 2 (Touvron et al., 2023b) for tabular and textual QA using the Step-wise Pipeline, which supports better for practical usage with cost and privacy concerns.
  • We conduct extensive experiments on three benchmarks, which validate the superiority of our TAT-LLM model over both conventional methods and very large-scale LMs like GPT-4, and the scaling law of LLMs on the tabular and textual QA task.
Figure 2. Comparison between a) End-to-end Pipeline and b) Step-wise Pipeline. c) Our TAT-LLM  language model is developed by fine-tuning LLaMA 2 (Touvron et al., 2023b) following the Step-wise Pipeline.
An example of QA with numerical reasoning over financial table-text data
Figure 2. Comparison between a) End-to-end Pipeline and b) Step-wise Pipeline. c) Our TAT-LLM  language model is developed by fine-tuning LLaMA 2 (Touvron et al., 2023b) following the Step-wise Pipeline.
2. Related Work
In this section, we review the related work on tabular and textual QA, large language model, and instruction tuning.
2.1. Tabular and Textual QA
Documents containing a hybrid of tabular and textual contents are very common on the Web (Zhu et al., 2021). The intelligent comprehension of these documents is often achieved with QA tasks. Early works (Chen et al., 2020b; Li et al., 2021, 2021; Chen et al., 2020a) focus on answer extraction from tabular and textual data. For example, HybridQA (Chen et al., 2020b) requires the model to extract a cell or a text span as the answer to a given question from Wikipedia tables and text. In recent two years, a growing body of works tackle tabular and textual QA by performing discrete reasoning as massive numerical information is usually included in the table or text, such as TAT-QA (Zhu et al., 2021), TAT-HQA (Li et al., 2022a), FinQA (Chen et al., 2021), MultiHiertt (Zhao et al., 2022), and TAT-DQA (Zhu et al., 2022). In this line of research, many supervised fine-tuning methods are proposed, such as TagOp (Zhu et al., 2021), DyRRen (Li et al., 2023), ELASTIC (Zhang and Moshfeghi, 2022), MVGE (Wei et al., 2023), RegHNT (Lei et al., 2022), SoarGraph (Zhu et al., 2023a) and UniRPG (Zhou et al., 2022). For example, TagOp (Zhu et al., 2021) first adopts sequence tagging to extract the supporting evidence from the table and text, and then applies pre-defined operators like addition, subtraction, and counting to infer the final answer. Harnessing the capabilities of advanced LLMs to tackle tabular and textual QA remains a relatively underexplored area, which is the main focus of our work.
2.2. Large Language Models (LLMs)
Recently, LLMs like ChatGPT (Ouyang et al., 2022a), GPT-4 (OpenAI, 2023) and PaLM (Chowdhery et al., 2022) have attracted tremendous research attention with their superb language understanding and generation abilities, bringing excellent performance to tasks of reading comprehension (Rogers et al., 2023), discrete reasoning(Wei et al., 2022c; Hendrycks et al., 2021; Cobbe et al., 2021b), etc. Since these general LLMs are usually extremely large and not accessible for most researchers, open-source smaller LLMs are proposed to allow researchers to customize their own LLMs for different tasks, such as LLaMA (Touvron et al., 2023a), Vicuna (Chiang et al., 2023), and Alpaca (Taori et al., 2023), with comparable performance to top general LLMs.
Existing research has attempted to instruct LLMs to perform table-based QA (Chen, 2023). Several works (Ye et al., 2023; Gemmell and Dalton, 2023; Jiang et al., 2023; Cheng et al., 2022) leverage general LLMs to solve the problem through a series of designed instructions or steps, which may include complex question decomposition, table evidence extraction, and SQL/code generation which is then executed via corresponding executors. However, these solutions do not focus on hybrid tabular and textual QA tasks, and rely on close-source general LLMs and carefully human-crafted instruction steps. Liu et al. (2023) focus on the multi-modal tabular, textual and image QA task, but still rely on instructing the general LLM to solve table-based questions. Our solution aims to enhance the tabular and textual QA ability by fine-tuning open-source LLMs.
2.3. Instruction Tuning
Instruction tuning is a very effective approach to boost the performance of pre-trained LLMs in various natural language tasks (Ouyang et al., 2022a; Zhang et al., 2023, 2023). It is an essential training stage for nearly all state-of-the-art LLMs (e.g., GPT-4 (OpenAI, 2023), Flan-T5 (Chung et al., 2022)), enabling pre-trained LLMs to understand the user query intent and follow user instructions. To perform instruction tuning, the LLM is trained on (input, output) pairs of natural language task data, where the input is usually a concatenation of task instruction, optional in-context examples and task input (Ouyang et al., 2022a). For general LLMs, instruction tuning usually aggregates massive data from a wide variety of natural language tasks (Chung et al., 2022) to develop well-round language abilities.
Apart from training general LLMs, some works also perform domain-specific instruction tuning to obtain LLMs for certain domains or tasks. For example, in financial domain, BloombergGPT (Wu et al., 2023) performs instruction tuning on a wide range of financial task data, largely outperforming general LLMs on financial tasks. Goat (Liu and Low, 2023) fine-tunes LLaMA on various arithmetic tasks, outperforming GPT-4 on some of these tasks. Model specialization (Fu et al., 2023) performs instruction tuning on small LMs to distill a certain task ability from general large LMs. Yang et al. (2023) distill table-based reasoning abilities from large LMs to smaller LMs. Similarly, our TAT-LLM focuses on financial tabular and textual QA tasks. To perform instruction tuning, techniques such as LoRA (Hu et al., 2022) and Qlora (Dettmers et al., 2023) are usually adopted to reduce the training cost.
3. Approach
In this section, we first introduce the Step-wise Pipeline for tabular and textual QA to highlight its merits, and then provide details of our TAT-LLM model specializing in discrete reasoning over hybrid tabular and textual data.
3.1. End-to-end Pipeline vs. Step-wise Pipeline
We consider such a problem: given a question and its relevant context, the model aims to derive the answer from the context, where discrete reasoning capabilities are usually required. In this work, the context refers to the documents with a hybrid of tabular and textual data.
3.1.1. End-to-end Pipeline
As shown in Figure 2 a), to solve the problem, the model that follows an End-to-end Pipeline takes the given question and context as input and directly outputs the answer. This straightforward approach provides little insight into how the answer is derived. It is akin to a black-box method where the internal workings remain largely concealed.
3.1.2. Step-wise Pipeline
We abstract this pipeline from previous multi-step methods such as (Zhu et al., 2021, 2023a; Li et al., 2023). As shown in Figure 2 b), this pipeline addresses the task with three steps: Extractor, Reasoner, and Executor. In particular, Extractor serves as an information extraction module that identifies the relevant snippets or segments of information to the question from the context; 2) Reasoner works as a “logic thinker” to generate the right logic rule or equation that leads to the right answer to the question; 3) Executor is responsible for deriving the final answer by performing the logic rule or calculating the equation.
Such a Step-wise Pipeline is able to produce a sequence of intermediate results, which enables us to boost each component within the pipeline, consequently leads to improved final performance. For example, in this work, we choose to strengthen the third step by adding an extra Executor, which brings noticeable performance gains as verified experimentally in Section 4.6.3.
3.1.3. Implementing Pipelines with LLMs
LLMs are shown to be able to follow user instructions in natural language to generate desired results (Ouyang et al., 2022b). It therefore has become a general consensus that LLMs can be trained to complete specific tasks as expected with appropriate instructions. In this work, our goal is to instruct LLMs to perform discrete reasoning over tabular and textual data following the Step-wise Pipeline. To achieve it, we carefully design a natural language instruction to include the three steps in the Step-wise Pipeline, i.e., Extractor, Reasoner, and Executor. After that, we generate the training data with publicly available tabular and textual QA datasets and train LLMs to answer the questions to follow the instructions. At below, we elaborate on how we develop our TAT-LLM language model following the Step-wise Pipeline, and provide a comprehensive comparison with the model trained following the End-to-end Pipeline in Section 4.6.2.
3.2. TAT-LLM
Here we introduce how we develop the TAT-LLM model.
3.2.1. Selection of the Language Model
We develop our TAT-LLM  by fine-tuning LLaMA 2  (Touvron et al., 2023b) on the training data transformed from publicly available expert-annotated tabular and textual QA datasets. LLaMA (Touvron et al., 2023a, b) is a series of open-source large language models that are trained on the large-scale publicly available corpus. It is commonly employed in the development of large language models with instruction tuning, such as Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), and Goat (Liu and Low, 2023). In this study, we choose the latest LLaMA 2 (Touvron et al., 2023b) as our base language model because it offers an extended maximum sequence length (i.e., $4,096$) compared to the previous version — LLaMA 1 (Touvron et al., 2023a), which often fails to accommodate the tabular and textual content of one instance from the existing QA datasets. Moreover, LLaMA 2 has demonstrated remarkable performance across various benchmarks.
Algorithm 1 : External Executor
Input $O_{1}$: the output of Extractor; $O_{2}$: the output of Reasoner; $O_{3}$: the output of Executor; $Q_{t}$: the predicted question type.
1:$answer\leftarrow O_{3}$
2:if $O_{2}$ is a valid arithmetic equation then
3:     $answer\leftarrow round(eval(O_{2}),4)$
4:else if “#” in $O_{2}$ then # multiple values are separated with “#”
5:     $arr\leftarrow O_{2}.split(``\#")$
6:     $answer\leftarrow len(arr)$
7:else if “¿” in $O_{2}$ or “¡” in $O_{2}$ then
8:     $answer\leftarrow eval(O_{2})$
9:else if $O_{2}$ = “N.A.” then
10:     if $Q_{t}$ = “Span” then
11:         $answer\leftarrow O_{1}$
12:     else if $Q_{t}$ = “Multiple Spans” then
13:         $arr\leftarrow O_{1}.split(``\#")$ spans are separated with “#”
14:         $answer\leftarrow arr$
15:     end if
16:end if
3.2.2. Construction of Training Data
To develop TAT-LLM, our fine-tuning is based on the training data we automatically transform from available tabular and textual QA datasets. For implementation in this work, we use three datasets, including FinQA (Chen et al., 2021), TAT-QA(Zhu et al., 2021) and TAT-DQA(Zhu et al., 2022), which are detailed in Section 4.1. We carefully design templates for each dataset to generate the training data for fine-tuning LLMs. Please refer to Appendix A for detailed information about their formats. As shown in Figure 2 c), each training instance is designed to be composed of five parts, including Instruction, Table, Text, Question, and Response.
Instruction: This part provides a detailed guide for what the task is and how to complete it. It outlines the steps that should be followed to derive the answer from the provided context. We adopt the Step-wise Pipeline described in above sections. 1) Extractor: We request the model to extract the relevant information from either the table or the accompanying text. 2) Reasoner: Based on the extracted information, we ask the model to derive an equation or a logic rule that is used to infer the answer. 3) Executor: We require the model to execute the equation or logic rule to arrive at the final answer. An illustration is given in Figure 2 c), which is the same for all the training instances. The instruction part helps the model to precisely understand the task and generate accurate and meaningful outputs.
Table: This part refers to the tabular data in the input context of this task. We compose this part using the tables from FinQA and TAT-QA datasets. As each table in the two datasets is stored in a Two-Dimensional (2D) array, we transform it into a markdown table. Note that we omit TAT-DQA dataset for constructing this part, considering the context given in TAT-DQA is the document pages from PDF documents, where the position and structure of the table are unknown.
Text: This part refers to the textual data in the input context. For FinQA and TAT-QA datasets, we compose this part with the annotated paragraphs associated with the table; all contents from the document pages in TAT-DQA dataset are regarded as textual data.
Question: This part refers to the question that is asked based on the input tabular and textual data. We directly put the question from the three datasets in this part. The goal of the task is to generate the right answer to this question.
Response: This part describes the model’s output based on the given table, text, and question. To better inspect the results in the intermediate steps and automatically evaluate the model output, we train the language model to format the output in a structured markdown table with two columns, i.e., step and output. Each row in the output table corresponds to one step defined in the instruction. In the end, the model draws its conclusion with a statement: ”The answer is: answer”, making its final response clear.
To train the model, the response part in the training set needs to be generated in advance. Existing tabular and textual QA datasets often provide the answer to the question and the derivation annotated by human experts. To generate the correct response for constructing our training instance, we first automatically identify the supporting evidence from the derivation of each question, such as the numbers and entities, which are used as the output of the first step. We then convert the derivation to a valid equation or logic rule as the output of the second step, making sure its execution result is consistent with the answer. We directly use the annotated answer as the output of the last step.
3.2.3. Training
After constructing the training set, we train our TAT-LLM model in various sizes, including 7B, 13B, and 70B, by fine-tuning LLaMA 2 (Touvron et al., 2023b) using Low-Rank Adaptation (LoRA) (Hu et al., 2022). More details are provided in Section 4.3.
3.2.4. External Executor
We observe that the trained model struggles in performing the execution of the mathematical equations or logic rules (e.g., counting, comparison) in the last step, according to experimental evaluations in Section 4.6.3. Such incompetence of the Executor has also been found in previous works (Luo et al., 2023; Yue et al., 2023; Liu and Low, 2023). To enhance the model’s accuracy of the final output, we propose to add an External Executor to our model, which takes the model’s intermediate outputs as input and performs the execution of the equations or logic rules to obtain the final answer. With this External Executor, the model is able to refine its output to a better one. With the output of the model formatted to a structured table, as aforementioned, we can easily access the intermediate results. The whole process is summarized in Algorithm 1. After obtaining the intermediate results, the External Executor is applied to refine the final answer instead of directly using the prediction of the model.
4. Experiments
We conduct extensive experiments to evaluate the effectiveness of our Step-wise Pipeline and TAT-LLM model. In addition to the comparisons with previous work, we also provide a comprehensive analysis to further verify our model design rationality, and reveal more properties of our method.
4.1. Datasets
We conduct the experiments using three publicly available tabular and textual QA datasets, i.e., FinQA (Chen et al., 2021), TAT-QA (Zhu et al., 2021) and TAT-DQA (Zhu et al., 2022), which are introduced at below. See Table 1 for the statistics of splits of each dataset.
FinQA (Chen et al., 2021) is an expert-annotated tabular and textual QA dataset in which the tables and text are sampled from financial reports. It focuses on discrete reasoning capabilities like addition, subtraction, multiplication, division, and numerical comparison.
TAT-QA (Zhu et al., 2021) is also built with tables and paragraphs extracted from financial reports. Most questions require discrete reasoning for generating the answers, and meanwhile there are also cases where the answers can be extracted directly from the tables or text. Its questions are classified into four different types: Span, Multiple Spans, Counting, and Arithmetic. In addition to the numerical value in the predicted answer, it also requires the model to predict the right scale of the answer, e.g., thousand, million.
TAT-DQA (Zhu et al., 2022) is an extension of the TAT-QA dataset, focusing on question answering over the original long financial statements with up to three pages, and the position and structure of the tables are unknown. This setting is believed to be more realistic and challenging than FinQA and TAT-QA datasets.
Table 1. Statistics of the train, validation and test splits for the financial tabular and textual QA datasets.
Dataset Train Set Dev Set Test Set
FinQA (Chen et al., 2021) 6,251 883 1,147
TAT-QA (Zhu et al., 2021) 13,215 1,668 1,669
TAT-DQA (Zhu et al., 2022) 13,251 1,645 1,662
4.2. Compared Models
We compare our TAT-LLM model with two kinds of models: fine-tuned models on the tabular and textual QA dataset, and LLMs in zero-shot setting. For fine-tuned models, we select the state-of-the-art supervised fine-tuned models for each dataset, which are listed separately in Table 2, Table 3 and Table 4. For LLMs, we utilize the state-of-the-art GPT3.5-Turbo (Brown et al., 2020) and GPT-4 (OpenAI, 2023). Besides, we also adopt powerful smaller LLMs that have garnered considerable research interest in the field, including Vicuna (7B) (Chiang et al., 2023) and LLaMA 2-Chat (7B and 70B) (Touvron et al., 2023b). In addition to these general LLMs, we also compare our models with LLMs specialized in math word problems, including MAmmoTH (70B) (Yue et al., 2023) and WizardMath (70B) (Luo et al., 2023), since our task involves numerical reasoning. All LLMs are tested in zero-shot setting, because financial tabular and textual data inputs tend to be lengthy, making it impractical to include additional in-context examples due to input length limits.
4.3. Implementation Details
Since the question type and the scale of the answer have been annotated in both TAT-QA and TAT-DQA datasets, we add two more steps in the instruction part of the input instance, in addition to the three steps in the proposed Step-wise Pipeline shown in Figure 2 c). One step is Question Type Predictor, with which we limit the model to select one value from the following question types: Span, Multiple Spans, Counting, and Arithmetic. For Span and Multiple Spans questions, we set the output of the Reasoner to “N.A.”. The other step is Scale Predictor, with which we restrict the model to choose one from the following values: Thousand, Million, Billion, Percent and None. For all the templates used to construct training instances of our  TAT-LLM model, please refer to Appendix A.
We train our TAT-LLM model by fine-tuning LLaMA 2 (Touvron et al., 2023b) using Low-Rank Adaptation (LoRA) (Hu et al., 2022) on one NVIDIA DGX-A100 with eight A100 GPUs. The quantization is 8 bit. We use Adam optimizer with a learning rate of $43e-$ 4 and warmup over the first $3\%$ steps to train. The maximum sequence length is $4,096$ and the maximum number of epochs is set to $3$. The batch size is set to $4$ and the gradient accumulation step is $10$.
For comparing with fine-tuned models, we take the results from their original papers respectively. For each LLM, we utilize three different templates to perform zero-shot inference to obtain prediction results. We select the best result as the reported result in Section 4.5. Please refer to Appendix B for details of the templates we use for zero-shot inference. We utilize the latest version [^1] Date: Sep 2023 GPT3.5-Turbo (Brown et al., 2020) and GPT-4 (OpenAI, 2023) via OpenAI APIs [^2] https://platform.openai.com/.. We set the temperature as 0, top p as 1.0, max token as $1,000$, and other parameters as default. We obtain the official trained checkpoint of Vicuna (Chiang et al., 2023), LLaMA 2-Chat (Touvron et al., 2023b), MAmmoTH (Yue et al., 2023) and WizardMath (Luo et al., 2023) from Hugginface [^3] https://huggingface.co/models. The inference is done on one NVIDIA DGX-A100 with eight A100 GPUs. The parameters num_beam and do_sample are 1 and false respectively.
4.4. Evaluation Metrics
We adopt the standard automatic evaluation metrics of FinQA, TAT-QA and TAT-DQA to evaluate the compared methods. For TAT-QA and TAT-DQA, we adopt the Exact Match (EM) and the numeracy-focused (macro-averaged) F1 score (Zhu et al., 2021, 2022). Both two metrics measure the overlap between a bag-of-words representation of the gold and predicted answers. The numeracy-focused F1 score is set to 0 unless the predicted number is exactly equal to the ground truth. We use EM for FinQA, which is the same as the metric of Execution Accuracy originally used in FinQA (Chen et al., 2021).
For zero-shot prediction using LLMs like GPT3.5-Turbo (Brown et al., 2020) and WizardMath (Luo et al., 2023), we omit the scale prediction and compare the predicted value with the ground truth value only, assuming the scale prediction is always correct. For our TAT-LLM model, since we instruct the model to output a structured table that is friendly to automatic evaluation, we take the scale into account.
4.5. Main Results
Table 2. Performance of our TAT-LLM model and compared models on the test set of FinQA. The best result is marked in bold and the second-best result is underlined. The number in red indicates the improvement over the second-best result.
Type Model EM
Human Expert Performance 91.16
Fine-tuned Longformer (Beltagy et al., 2020) 21.90
NeRd (Chen et al., 2019) 48.57
[…]
Zero-shot Vicuna (7B) (Chiang et al., 2023) 10.11
LLaMA 2-Chat (7B)  (Touvron et al., 2023b) 15.43
LLaMA 2-Chat (70B)  (Touvron et al., 2023b) 32.17
MAmmoTH (70B) (Yue et al., 2023) 36.09
WizardMath (70B) (Luo et al., 2023) 47.25
GPT-3.5-Turbo (Brown et al., 2020) 58.00
GPT-4 (OpenAI, 2023) 63.91
Ours TAT-LLM (7B) 64.60 (+0.69)
We first compare the performance of our TAT-LLM  with compared methods on each tabular and textual QA dataset, i.e., FinQA, TAT-QA and TAT-DQA, respectively. The experimental results are summarized in Table 2 for FinQA, Table 3 for TAT-QA and Table 4 for TAT-DQA. From the tables, we make the following observations. 1) Our TAT-LLM (7B) significantly outperforms all other models on all three datasets, including the previous best fine-tuned models and the state-of-the-art GPT-4 (OpenAI, 2023). In particular, our TAT-LLM reaches $64.60\%$, $74.56\%$ and $69.45\%$ in terms of EM on the test set of FinQA, TAT-QA and TAT-DQA respectively, i.e. an increase of $0.69$, $2.64$ and $4.99$ points compared to GPT-4. These results well demonstrate the great effectiveness of our model. The results also confirm the rousing potential of specializing smaller language models, like our TAT-LLM (7B) model, for specific tasks to yield better performance than very large-scale models like GPT-4. 2) GPT-4 with our Step-wise Pipeline surpasses the best fine-tuned models on all datasets, showing the rationality of such a pipeline. 3) However, the performance of GPT-4 and our TAT-LLM  obviously lags behind that of human experts, showing that this task is still challenging. 4) The strong general LLM GPT3.5-Turbo and LLaMA 2-Chat (70B) underperform the best fine-tuned models, evidencing that supervised fine-tuning is still essential for achieving advanced performance on table-text QA datasets. 5) The LLMs specialized in math word problems, i.e., WizardMath (Luo et al., 2023) and MAmmoTH (Yue et al., 2023), largely underperform GPT3.5-Turbo, GPT-4 and our TAT-LLM model, showing that current numerically-enhanced LLMs still struggle in discrete reasoning over tabular and textual QA. This again speaks for the considerable value of specializing models for different tasks.
Table 3. Performance of our TAT-LLM model and compared models on the test set of TAT-QA. Best results are marked in bold and second-best results are underlined. Numbers in red indicate the improvement over the second-best results.
Type Model EM F1
Human Expert Performance 84.1 90.8
Fine-tuned TagOp (Zhu et al., 2021) 50.10 58.00
TeaBReaC (Trivedi et al., 2022) 55.80 63.80
KIQA (Nararatwong et al., 2022b) 58.20 67.40
FinMath (Li et al., 2022b) 58.30 68.20
GANO (Nararatwong et al., 2022a) 61.90 72.10
MHST (Zhu et al., 2022) 63.60 72.70
UniPCQA (Deng et al., 2022) 63.90 72.20
SoarGraph (Zhu et al., 2023b) 65.40 75.30
UniRPG (Zhou et al., 2022) 67.20 76.00
RegHNT (Lei et al., 2022) 70.30 77.90
MVGE (Wei et al., 2023) 70.90 79.10
Zero-shot Vicuna (7B) (Chiang et al., 2023) 32.53 40.97
LLaMA 2-Chat (7B) (Touvron et al., 2023b) 37.16 45.37
MAmmoTH (70B) (Yue et al., 2023) 38.97 46.51
WizardMath (70B) (Luo et al., 2023) 39.63 45.28
LLaMA 2-Chat (70B) (Touvron et al., 2023b) 45.94 53.80
GPT-3.5-Turbo (Brown et al., 2020) 59.47 68.11
GPT-4 (OpenAI, 2023) 71.92 79.71
Ours TAT-LLM (7B) 74.56 (+2.64) 82.88 (+3.17)
4.6. Model Analysis
4.6.1. Fine-tuning with Combined Data.
[…]
Table 4. Performance of our TAT-LLM  model and compared models on the test set of TAT-DQA. Best results are marked in bold and second-best results are underlined. Numbers in red indicate the improvement over the second-best results.
Type Model EM F1
Human Expert Performance 84.1 90.8
Fine-tuned NumNet+ V2 (Ran et al., 2019) 30.60 40.10
TagOp (Zhu et al., 2021) 33.70 42.50
MHST (Zhu et al., 2022) 41.50 50.70
Doc2SoarGraph (Zhu et al., 2023b) 59.20 67.60
Zero-shot Vicuna (7B) (Chiang et al., 2023) 28.44 36.72
LLaMA 2-Chat (7B)  (Touvron et al., 2023b) 34.52 42.32
MAmmoTH (70B) (Yue et al., 2023) 35.42 42.82
WizardMath (70B) (Luo et al., 2023) 36.44 41.55
LLaMA 2-Chat (70B)  (Touvron et al., 2023b) 41.91 49.74
GPT-3.5-Turbo (Brown et al., 2020) 52.74 61.40
GPT-4 (OpenAI, 2023) 64.46 72.20
Ours TAT-LLM (7B) 69.45 (+4.99) 77.75 (+5.55)
Table 5. Performance of the TAT-LLM model on the test sets, which is trained using a combination of all three training sets. Best results are marked in bold. Numbers in red indicate the improvement compared with GPT-4.
Model FinQA TAT-QA TAT-DQA
EM EM F1 EM F1
GPT-3.5-Turbo 58.00 59.47 68.11 52.74 61.40
GPT-4 63.91 71.92 79.71 64.46 72.20
Fine-tuned with respective training set:
TAT-LLM (7B) (+0.69) (+2.64) (+3.17) (+4.99) (+5.55)
64.60 74.56 82.88 69.45 77.75
Fine-tuned with a combination of training sets:
[…]
65.13 76.49 85.13 71.38 80.24
[…]
71.93 77.51 85.95 72.22 80.56
[…]
76.81 81.42 88.49 76.55 83.90
4.6.2. Different Fine-tuning Strategies.
[…]
Model FinQA TAT-QA TAT-DQA
EM EM F1 EM F1
[…]
  w/o External Executor 48.47 58.69 67.21 54.84 63.68
  w External Executor 65.13 76.49 85.13 71.38 80.24
  gains (+16.66) (+17.80) (+17.92) (+16.54) (+16.56)
[…]
  w/o External Executor 60.05 62.60 70.73 59.95 68.61
  w External Executor 71.75 76.79 85.05 71.86 80.50
  gains (+11.70) (+14.19) (+14.32) (+11.91) (+11.89)
[…]
  w/o External Executor 70.10 76.61 83.55 71.74 78.99
  w External Executor 76.81 81.42 88.49 76.55 83.90
  gains (+6.71) (+4.81) (+4.94) (+4.81) (+4.91)
GPT-4 63.91 71.92 79.71 64.46 72.20
To verify the effect of our fine-tuning strategy, i.e. Step-wise Pipeline plus External Executor, we compare it with separately applying End-to-end Pipeline or Step-wise Pipeline on LLaMA 2 (7B) models. The results are summarized in Figure 3. Firstly, we can see our TAT-LLM model clearly outperforms the two variants following the End-to-end Pipeline and the Step-wise Pipeline on all tabular and textual QA datasets for both evaluation metrics, demonstrating the effectiveness of our fine-tuning strategy. Besides, we find that the Step-wise Pipeline also outperforms the End-to-end Pipeline in all situations, showing the rationality of incorporating intermediate reasoning steps, which is in line with previous research on chain-of-thoughts reasoning (Wei et al., 2022c; Chung et al., 2022). Moreover, it is observed that the End-to-end Pipeline achieves notably poor performance on FinQA compared to that on TAT-QA and TAT-DQA. This is probably because FinQA requires more complex discrete reasoning steps than TAT-QA and TAT-DQA, and omitting such steps in fining-tuning would largely degrade model performance. Finally, we find a significant performance increase from the Step-wise Pipeline to TAT-LLM. This is due to the incorporation of the reliable External Executor which largely guarantees the reasoning accuracy of the answer. In the next section, we perform detailed analysis on the effectiveness of the External Executor.
Figure 3. Comparison on different fine-tuning strategies.
An example of QA with numerical reasoning over financial tabular and textual data
Figure 3. Comparison on different fine-tuning strategies.
4.6.3. Effect of External Executor w.r.t. Different Model Sizes.
[…]
Figure 4. Performance comparison in terms of EM between TAT-LLM (7B) and LLaMA 2-Chat (7B) for different question types on TAT-QA. Numbers in red denote the relative improvement from LLaMA 2-Chat (7B) to TAT-LLM (7B).
An example of QA with numerical reasoning over financial tabular and textual data
Figure 4. Performance comparison in terms of EM between TAT-LLM (7B) and LLaMA 2-Chat (7B) for different question types on TAT-QA. Numbers in red denote the relative improvement from LLaMA 2-Chat (7B) to TAT-LLM (7B).
Figure 5. Performance comparison in terms of EM between TAT-LLM (7B) and LLaMA 2-Chat (7B) for different question types on TAT-DQA. Numbers in red denote the relative improvement from LLaMA 2-Chat (7B) to TAT-LLM (7B).
4.6.4. Improvements w.r.t. Different Question Types.
Here we analyze the improvements achieved by TAT-LLM model regarding different question types of TAT-QA and TAT-DQA datasets. We compare the performance of TAT-LLM (7B) and LLaMA 2-Chat (7B) for each question type within TAT-QA and TAT-DQA. The results are summarized in Figure 4 and Figure 5 for TAT-QA and TAT-DQA, respectively. We can observe substantial performance improvements for the Arithmetic questions, which are over two-fold, for both datasets. The larger performance gains are attributed to the effective fine-tuning strategy of TAT-LLM  as well as the reliable External Executor that ensures accurate execution of the equations and logic rules. Large performance improvements are also found for the Counting questions, namely 109.1% for TAT-QA and 53.8% for TAT-DQA. This is probably due to executing more correct counting operations. For the Span and Multiple Spans types, relatively smaller improvements are observed, which are still over 30%. This is potentially because these questions involve simpler or no discrete reasoning, mainly numerical comparison.
Table 7. Examples of errors on the test set of FinQA and corresponding percentages in each step. Q, G and P denote question, ground truth and prediction.
Step Error(%) Example
Extractor Wrong Evidence (48%)
Q: What is the percentage change in cash flow hedges in 2011 compared to 2010?
G: 153.7, 139.9
P: 153.7, 375.0
Missed Evidence (15%)
Q: What was the total impairment costs recorded from 2003 to 2005 in millions?
G: 0.6, 0.5, 4.7
P: 0.6, 0.5
Redundant Evidence (8%)
Q: What was the average number of shares issued to employees from 2013 to 2015?
G: 439000, 411636
P: 439000, 411636, 556000
Reasoner Wrong Operators (19%)
Q: what was the percent of the change in the stock price performance for hum from 2010 to 2011?
G: (201 - 125) / 125
P: 201 - 125
Wrong Values (10%)
Q: What was the average cash flow from 2004 to 2006?
G: (950.4 + 957.4 + 769.1) / 3
P: (957.4 + 957.4 + 769.1) / 3
4.6.5. Error Analysis
To further diagnose our TAT-LLM model, we randomly sample $100$ error instances of TAT-LLM (7B) from the test set of FinQA and analyze the reasons. Since we adopt the External Executor for reliable execution, the errors only occur to the Extractor and Reasoner. As shown in Table 7, we list all kinds of errors with examples and their corresponding percentages. We can observe that most of the errors come from the Extractor, where the Wrong Evidence takes $48\%$ of the total errors. This shows that our TAT-LLM (7B) model still faces challenges in precisely interpreting the meaning of the values, sometimes due to the unique terminology. Training the model with more tabular and textual data in this domain might be a possible approach to enhancing its evidence extraction ability. Additionally, the Wrong Operator and the Wrong Values may be partially caused by the randomness of the LLM as the generation length increases, which fails to generate the full equation or wrongly copies the extracted evidence.
5. Conclusion
In this work, we first abstract a Step-wise Pipeline for tabular and textual QA based on the previous multi-step methods. Following this pipeline, we develop the TAT-LLM model by specializing smaller language models (i.e., LLaMA 2) in discrete reasoning over tabular and textual data. We validate its effectiveness with extensive experiments and analyses, which show that our TAT-LLM model can outperform both conventional methods and very large-scale LMs like GPT-4 on this task. Our work well demonstrates that specialized models might be a promising direction towards more advanced models on specific tasks that can compete with human expertise.
Appendix A Templates for Fine-tuning
To prepare data for training our TAT-LLM model, the below templates are applied to construct the training instances following the Step-wise Pipeline on the three tabular and textual QA datasets, i.e., Table A1 for FinQA, Table A2 for TAT-QA and Table A3 for TAT-DQA, respectively.
For comparison, we also present the templates we used to prepare the training instances following the End-to-end Pipeline in Table A4 for FinQA, Table A5 for TAT-QA, and Table A6 for TAT-DQA.
Table 1. The template for constructing training instances following Step-wise Pipeline with FinQA dataset.
Below is an instruction that describes a question answering task in the finance domain, paired with an input table and its relevant text that provide further context. The given question is relevant to the table and text. Generate an appropriate answer to the given question.
### Instruction:
Given a table and a list of texts in the following, what is the answer to the question? Please complete the task in three steps:
1. In the first step, extract the relevant numerical values from the provided table or texts. Store these in the variable ‘{evidence}‘. If there are multiple values, separate them using the ’#’ symbol.
2. In the second step, generate an equation using the extracted numerical values. Store this equation in the variable ‘{equation}‘.
3. In the third step, calculate the answer based on the equation and store it in the variable ‘{answer}‘.
Please organize the results in the following table:
— step — output —
— 1 — {evidence} —
— 2 — {equation} —
— 3 — {answer} —
Finally, present the calculated answer in the format: ”The answer is: {answer}”
### Table
{table}
### Text
{text}
### Question
{question}
### Response
—step — output—
—1 — {gold_evidence} —
—2 — {gold_equation} —
—3 — {gold_answer} —
The answer is: {gold_answer}
Table 2. The template for constructing training instances following Step-wise Pipeline with TAT-QA dataset.
Below is an instruction that describes a question answering task in the finance domain, paired with an input table and its relevant text that provide further context. The given question is relevant to the table and text. Generate an appropriate answer to the given question.
### Instruction
Given a table and a list of texts in the following, answer the question posed using the following five-step process:
1. Step 1: Predict the type of question being asked. Store this prediction in the variable ‘{question_type}‘. The value of ‘{question_type}‘ can be one of the following:‘Single span‘, ‘Multiple spans‘, ‘Count‘, or ‘Arithmetic‘.
2. Step 2: Extract the relevant strings or numerical values from the provided table or texts. Store these pieces of evidence in the variable ‘{evidence}‘. If there are multiple pieces of evidence, separate them using the ’#’ symbol.
3. Step 3: if the ‘{question_type}‘ is ‘Arithmetic‘, formulate an equation using values stored in ‘{evidence}‘. Store this equation in the variable ‘{equation}‘. For all other question types, set the value of {equation} to ’N.A.’.
4. Step 4: Predict or calculate the answer based on the question type, evidence and equation. Store it in the variable ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol.
5. Step 5: If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’.
Please organize the results in the following table:
— step — output —
— 1 — {question_type} —
— 2 — {evidence} —
— 3 — {equation} —
— 4 — {answer} —
— 5 — {scale} —
Finally, present the final answer in the format: ”The answer is: {answer} #### and its corresponding scale is: {scale}”
### Table
{table}
### Text
{text}
### Question
{question}
### Response
— step — output —
— 1 — {gold_question_type} —
— 2 — {gold_evidence} —
— 3 — {gold_equation} —
— 4 — {gold_answer} —
— 5 — {gold_scale} —
The answer is: {gold_answer} #### and its corresponding scale is: {gold_scale}
Table 3. The template for constructing training instances following Step-wise Pipeline with TAT-DQA dataset.
Below is an instruction that describes a question answering task in the finance domain, paired with an input document that has one or multiple pages that provide further context. The given question is relevant to the document. Generate an appropriate answer to the given question.
### Instruction
Given a document that has one or multiple pages in the following, answer the question posed using the following five-step process:
1. Step 1: Predict the type of question being asked. Store this prediction in the variable ‘{question_type}‘. The value of ‘{question_type}‘ can be one of the following:‘Single span‘, ‘Multiple spans‘, ‘Count‘, or ‘Arithmetic‘.
2. Step 2: Extract the relevant strings or numerical values from the provided document. Store these pieces of evidence in the variable ‘{evidence}‘. If there are multiple pieces of evidence, separate them using the ’#’ symbol.
3. Step 3: if the ‘{question_type}‘ is ‘Arithmetic‘, formulate an equation using values stored in ‘{evidence}‘. Store this equation in the variable ‘{equation}‘. For all other question types, set the value of {equation} to ’N.A.’.
4. Step 4: Predict or calculate the answer based on the question type, evidence and equation. Store it in the variable ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol.
5. Step 5: If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’.
Please organize the results in the following table:
— step — output —
— 1 — {question_type} —
— 2 — {evidence} —
— 3 — {equation} —
— 4 — {answer} —
— 5 — {scale} —
Finally, present the final answer in the format: ”The answer is: {answer} #### and its corresponding scale is: {scale}”
### Text
{pages}
### Question
{question}
### Response
— step — output —
— 1 — {gold_question_type} —
— 2 — {gold_evidence} —
— 3 — {gold_equation} —
— 4 — {gold_answer} —
— 5 — {gold_scale} —
The answer is: {gold_answer} #### and its corresponding scale is: {gold_scale}
Table 4. The template for constructing training instances following End-to-end Pipeline with FinQA dataset
Below is an instruction that describes a question answering task in the finance domain, paired with an input table and its relevant text that provide further context. The given question is relevant to the table and text. Generate an appropriate answer to the given question.
### Instruction
Given a table and a list of texts in the following, what is the answer to the question? Please output the answer in the format of ”The answer is:”.
### Table
{table}
### Text
{text}
### Question
{question}
### Response
The answer is: {answer}
Table 5. The template for constructing training instances following End-to-end Pipeline with TAT-QA dataset
Below is an instruction that describes a question answering task in the finance domain, paired with an input table and its relevant text that provide further context. The given question is relevant to the table and text. Generate an appropriate answer to the given question.
### Instruction
Given a table and a list of texts in the following, what is the answer to the question? Please predict the answer and store it in a variable named ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol. If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’. Finally, present the final answer in the format of ”The answer is: {answer} #### and its corresponding scale is: {scale}”
### Table
{table}
### Text
{text}
### Question
{question}
### Response
The answer is: {gold_answer} #### and its corresponding scale is: {gold_scale}
Table 6. The template for constructing training instances following End-to-end Pipeline with TAT-DQA dataset
Below is an instruction that describes a question answering task in the finance domain, paired with an input document that has one or multiple pages that provide further context. The given question is relevant to the document. Generate an appropriate answer to the given question.
### Instruction
Given a document that has one or multiple pages in the following, what is the answer to the question? Please predict the answer and store it in a variable named ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol. If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’. Finally, present the final answer in the format of ”The answer is: {answer} #### and its corresponding scale is: {scale}”
### Document
{pages}
### Question
{question}
### Response
The answer is: {gold_answer} #### and its corresponding scale is: {gold_scale}
Appendix B Templates for Zero-shot Inference
For zero-shot inference with the baseline LLMs such as GPT-3.5-Turbo (Brown et al., 2020) and GPT-4 (OpenAI, 2023), we utilize three different templates to build the instances and feed them to the LLMs to get the predictions. We report the best results in Section 4.5. The following are details of the templates we used. 1) End-to-end Pipeline template, i.e., Table A7 for FinQA and TAT-QA, and Table A8 for TAT-DQA; 2) Chain-of-Thought template, which adds ”Let’s complete this task step by step” to the End-to-end Pipeline template; 3) Step-wise Pipeline template that is similar to the templates for fine-tuning, with the part of “Response” removed.
Table 7. The template for zero-shot inference using the baseline LLMs on FinQA and TAT-QA datasets.
Below is an instruction that describes a question answering task in the finance domain, paired with an input table and its relevant text that provide further context. The given question is relevant to the table and text. Generate an appropriate answer to the given question.
### Instruction
Given a table and a list of texts in the following, what is the answer to the question? (Let’s complete this task step by step.) Please output the answer in the format of ”The answer is:”.
### Table
{table}
### Text
{text}
### Question
{question}
### Response
Table 8. The template for zero-shot inference using the baseline LLMs on TAT-DQA datasets.
Below is an instruction that describes a question answering task in the finance domain, paired with an input document that has one or multiple pages that provide further context. The given question is relevant to the document. Generate an appropriate answer to the given question.
### Instruction
Given a document with one or multiple pages in the following, what is the answer to the question? (Let’s complete this task step by step.) Please output the answer in the format of ”The answer is:”.
### Text
{pages}
### Question
{question}
### Response